What is the identity of substance X if 0.380 mol of X weighs 17.5 g? ОА. N2O4 O B. NO3 ос. No2 ор. N20
Answer #1


Use:
Mol of X = mass of X / Molar mass of X
0.380 mol = 17.5 g / Molar mass of X
Molar mass of X = 46.05 g/mol

A)

Molar mass of N2O4,
MM = 2*MM(N) + 4*MM(O)
= 2*14.01 + 4*16.0
= 92.02 g/mol

B)
Molar mass of NO3,
MM = 1*MM(N) + 3*MM(O)
= 1*14.01 + 3*16.0
= 62.01 g/mol

C)
Molar mass of NO2,
MM = 1*MM(N) + 2*MM(O)
= 1*14.01 + 2*16.0
= 46.01 g/mol

D)
Molar mass of N2O,
MM = 2*MM(N) + 1*MM(O)
= 2*14.01 + 1*16.0
= 44.02 g/mol

Answer: C. NO2